# Various Properties of Various Ultrafilters, Various Graph Width Parameters, and Various Connectivity Systems

### v6

@@ 6.2.6 Various Ultrafilters on a Connectivity System
Ultrafilters are of interest due to their practical applications and mathematical properties, prompting the proposal of various similar concepts across different fields. Going forward, we aim to define concepts such as Partition-filter [377], uniform-ultrafilters [557], weak normal-ultrafilter [433], gw-ultrafilters [475], subuniform ultrafilters [556], Fuzzy-ultrafilter [646],Regular-ultrafilters [502], Good-ultrafilters [166], Semi-ultrafilters [520], Linear-ultrafilters [58, 378], OK-ultrafilters [207], complete-ultrafilters [594], Ramsey-ultrafilters [589], and Selective-ultrafilters [589] within the frameworks of Undirected Graph, Directed Graph, Infinite Graph, Connectivity Systems and Infinite Connectivity Systems. Additionally, we plan to explore their relationships with graph width parameters.
@@ A.1 Tree-width
Tree-width is a measure of how tree-like a graph is, defined by its minimum tree decomposition width [613, 84, 608, 609, 624, 396, 612]. The definitions of Tree-width on a graph are provided below.
@@ A.1.1 Hierarchy for treewidth
  • Let $G$ be a connected graph with treewidth $tw(G)$ and gonality $dgon(G)$. Then $tw(G)\leq dgon(G)$ [678].
  • If $G$ has tree-width $k$, the dimension of $G$ is at most $k$ [85].
@@
  • If $G$ has tree-width $k-1$, $G$ has haven $k$ [625].
  • If $G$ has tree-width $k$, $G$ has no screen of thickness at least $k+2$ [85].
@@
  • The stack-number $sn(G)\leq tw(G)+1$ [328].
@@
  • If a graph $G$ has bounded tree-width, then it also has bounded Bramble Number (bn), Tangle Number, Lexicographic Tree Product Number (LTP), Linkedness Number (link), Cartesian Tree Product Number (ctp), and Well-Linked Number (wl). The converse is also true [384].
@@
Proof A.4.
@@ A.1.2 Related concepts for tree-width
Related concepts include maximum order of a grid minor, maximum order of a grid-like-minor, Hadwiger number, fractional Hadwiger number, domino treewidth [88, 86], and the following concepts [358].
@@
  • Linear Local Tree-width: A linear variation of Local Tree-width.
  • Simple Tree-width:
@@
  • Circuit Tree-width [605, 113]:
@@
  • Connected Tree-width [198, 554]: The minimum width of a tree-decomposition whose parts induce connected subgraphs.
@@ A.1.3 Application aspect for tree-width
We consider about application aspect for tree-width. Due to the versatility of tree structures in handling various types of data, tree-width has numerous applications across different fields.
@@ A.2 Path-width
Path-width [453, 454, 372, 442] quantifies how similar a graph is to a path by determining the minimum width of its path decomposition. An example of an obstruction to low path-width is blockage [73]. Path-width is notably applied in VLSI design [571]. Path-width is also called interval-width [459, 265].
As an example, the definition of Path-width in the context of matroids is provided. For the definition of a matroid, please refer to Definition A.24 and related references.
@@ A.2.1 Hierarchy for path-width
Proof A.7.
@@ A.2.2 Related concepts for path-width
  • Proper-path-width [656, 653, 657, 655, 654, 532, 531]: Path-width that relates to mixed search game (A pursuit-evasion game on graphs where a searcher combines edge and node searches to capture a hidden intruder).
@@
  • Connected path-width [179]: Linear layout parameter of Connected tree-width. Connected pathwidth of any graph $G$ is at most $12\times pw(G)+$ 1 [180].
@@ A.5 Linear MIM-width
Linear MIM-width [412, 413] is the linear restriction of MIM-width. If a graph $G$ has bounded linear-MIM-width, then it also has bounded MIM-width (trivially).
A.5.1 Related concepts for Linear-MIM-width
@@ A.8.1 Hierarchy for Band-width
  • If a graph $G$ has bounded band-width, then it also has bounded maximum degree, $c$-closure [467], acyclic chromatic number, and $h$-index [240] [668]  [^10] The $h$-index of a graph is the maximum integer $h$ such that the graph has at least $h$ vertices, each with a degree of at least $h$(cf.[241, 74])..
@@
  • The slope number $\text{sn}(G)$ satisfies: $\text{sn}(G)\leq\frac{1}{2}\text{bw}(G)\cdot(\text{bw}(G)+1)+1.$[218]
Proof A.18.
@@ A.9 Boundary-width
Vertex-boundary-width [576] measures the minimum boundary of vertices in a discrete isoperimetric problem. Note that the discrete isoperimetric problem seeks to determine the minimum "boundary" size of a set of vertices in a graph, given a fixed number of vertices in that set, often relating to optimizing shapes in grid-like structures or networks.
Edge-boundary-width [576] measures the minimum boundary of edges in a discrete isoperimetric problem.
@@ A.10 Carving-width
A.10.1 Related concept for Carving-width
@@
A.10.2 Application for carving-width
@@ A.11 Branch-width
Branch-width [330, 577, 613, 95] measures the smallest width of a branch decomposition of a graph. The width is the minimum number of edges connecting subgraph pairs. This graph width parameter is known to be extendable to matroids and connectivity systems. Examples of obstructions are Tangle [330, 393], Profile [192, 193, 238, 190, 199], $k$-block [231, 244, 464], Ultrafilter [304], Maximal ideal [713], loose tangle [577], Quasi-Ultrafilter [293], Weak-Ultrafilter [288], tangle kit[407], Bramble[516, 24], (k,m)-obstacle [269]. Related concepts include sphere-cut width [75] and Rooted Branch-width[418].

### v8

@@ 6.2.6 Various Ultrafilters on a Connectivity System
Ultrafilters are of interest due to their practical applications and mathematical properties, prompting the proposal of various similar concepts across different fields. Going forward, we aim to define concepts such as Partition-filter [629], uniform-ultrafilters [938], weak normal-ultrafilter [725], gw-ultrafilters [795], subuniform ultrafilters [936], Fuzzy-ultrafilter [1116], Regular-ultrafilters [836], Good-ultrafilters [297], Semi-ultrafilters [865], Dfin-ultrafilter [800], $I$-Ultrafilter [456], Open-Ultrafilter [245], Closed-Ultrafilter [121], Linear-ultrafilters [125, 630], OK-ultrafilters [371], Club-Ultrafilter [986], neutrosophic-ultrafilters [1048, 1241], complete-ultrafilters [1000], Ramsey-ultrafilters [985], and Selective-ultrafilters [985] within the frameworks of Undirected Graph, Directed Graph, Infinite Graph, Connectivity Systems and Infinite Connectivity Systems. Additionally, we plan to explore their relationships with graph width parameters.
@@ A.1 Tree-width
Tree-width is a measure of how tree-like a graph is, defined by its minimum tree decomposition width [1038, 164, 1033, 1034, 1071, 665, 1037]. Many algorithms that are NP-hard for general graphs, become easier when the treewidth is bounded by a constant. The definitions of Tree-width on a graph are provided below.
@@
Example A.3.
(cf.[516]) Consider a complete graph $K_{n}$ with $n$ vertices. Every pair of vertices is connected by an edge, so $E(K_{n})=\{(v_{i},v_{j})\mid 1\leq i<j\leq n\}$.
  • Nodes and Bags:
    • A tree-decomposition of $K_{n}$ can consist of a single node tree $T$ with one bag containing all vertices:
      $W=\{v_{1},v_{2},\dots,v_{n}\}.$
  • Tree $T$:
    • $T$ has only one node associated with the bag $W=V(K_{n})$.
  • Properties Verification:
    • Vertex Containment (T1):
      $\bigcup_{t\in V(T)}W=V(K_{n}).$
    • Edge Containment (T2):
        For each edge $(v_{i},v_{j})\in E(K_{n})$, both $v_{i}$ and $v_{j}$ are in the bag $W$.
    • Vertex Connectivity (T3):
        There is only one bag, so this condition is trivially satisfied.
  • Width Calculation:
    • The size of the bag $W$ is $|W|=n$.
    • The width of this tree-decomposition is $n-1$.
    • Since no tree-decomposition can have a smaller width, the Tree-width of $K_{n}$ is:
      $\text{TreeWidth}(K_{n})=n-1.$
@@ A.1.1 Hierarchy for treewidth
  • Let $G$ be a connected graph with treewidth $tw(G)$ and gonality $dgon(G)$. Then $tw(G)\leq dgon(G)$ [1170].  [^15] The gonality of a graph is the minimum number of chips needed to reach any vertex using chip-firing moves without causing debt at other vertices[291].
  • If $G$ has tree-width $k$, the dimension of $G$ is at most $k$ [165].  [^16] In graph theory, the dimension of a graph refers to the minimum size of cliques formed during vertex elimination, minimized over all possible perfect elimination orderings.
@@
  • If $G$ has tree-width $k-1$, $G$ has haven $k$ [1072].  [^17] A haven in a graph is a function that maps each subset of vertices of size at most $k$ to a connected component, ensuring that any two components ”touch” either by intersection or adjacency.
  • If $G$ has tree-width $k$, $G$ has no screen of thickness at least $k+2$ [165].  [^18] Screen in a graph is a collection of connected subsets of vertices where each pair of subsets either intersects or contains adjacent vertices, ensuring they ”touch.”
@@
  • The stack-number $sn(G)\leq tw(G)+1$ (stack-number is also known as book thickness or page-number) [560].
@@
  • If a graph $G$ has bounded Tree-width, then it also has bounded tree-independence number [116].
@@
  • If a graph $G$ has bounded tree-width, then it also has bounded Bramble Number (bn), Tangle Number, Lexicographic Tree Product Number (LTP) [^20] The lexicographic tree product number of a graph $G$, denoted by $ltp(G)$, is the smallest integer $k$ such that $G$ is a minor of $T\cdot K_{k}$, where $T$ is a tree and $K_{k}$ is the complete graph on $k$ vertices[642]., Linkedness Number (link)[1030], Cartesian Tree Product Number (ctp)[1168, 320]  [^21] The Cartesian Tree Product Number $\text{ctp}(G)$ is the smallest integer $k$ such that a graph $G$ is a minor of $T^{(k)}$, where $T^{(k)}$ is the Cartesian product of a tree $T$ and $K_{k}$[642]., and Well-Linked Number (wl)[1030]. The converse is also true [642].
@@
  • For any graph $G$, $\text{Cubicity}(G)\leq(\text{treewidth}(G)+2)\left\lfloor\log_{2}n\right\rfloor$ [260].
@@
  • Let $G$ be a graph. Then surrounding cop number is less than or equal to tree-width plus one[233, 1067].
  • For every connected graph, periodic cop number is less than or equal to tree-width plus one[1088].
  • For every connected graph, 00-domination cop number equals to tree-width plus one[467].
  • For every graph, $\infty$-colouring number equals to tree-width plus one[1167].  [^28] Also, for every graph, weak $\infty$-colouring number equals to tree-depth[1167].
[…]
  • Graphs that have a treewidth of exactly $k$ are referred to as $k$-trees, while those with a treewidth of at most $k$ are known as partial $k$-trees [162].
  • If $tw(G)\leq k$, then maximum interval-number is $k+1$[777], track-number is $k+1$[777], and bend-number is $22k-$ 2[654].  [^29] The bend-number of a graph is the minimum $k$ such that it can be represented by grid paths with at most $k$ bends each, where two vertices are adjacent if and only if their corresponding paths share a grid edge[654, 653].
  • Block number is less than or equal to treewidth + 1[1204].
  • If $G$ has treewidth $k$, it is known to have a balanced separator of size $k+1$[153].
  • Let $G$ be a graph on $n$ vertices. Then $\text{TW}(G)\leq k$ if and only if there exists a $k$-achievable set $S$ of size $n-k-1$ [850].
  • Let $G$ be a graph and $H$ a minor of $G$. Then $\text{tw}(H)\leq\text{tw}(G)$ [311].
  • Let $G$ be a graph. Then $\textit{treewidth}(G)\leq\textit{catwidth}(G)\leq\textit{pathwidth}(G)$[1140].  [^30] Catwidth is the minimum tree-decomposition width of a graph $G$ when the pathwidth of the decomposition tree $T$ is restricted to 1[1140].
  • Let $G$ be a graph. Then comparable box $\textit{comparable box dimension}\leq\textit{treewidth}(G)+1$[400].
  • Every graph $G$ has a 2-blocking partition with width at most $1350(\text{tw}(G)+1)(\Delta(G)^{2}.$ ($\Delta$: maximum degree) [359]  [^31] An $\ell$-blocking partition of a graph $G$ is a partition of the vertex set $V(G)$ into connected sets such that every path of length greater than $\ell$ in $G$ contains at least two vertices in one part of the partition[359].
  • For a connected graph $G$, we have the following inequality:
    $\text{tw}(G)<\Delta(G)\left(\text{spanning tree congestion}(G)+1\right),$
    where $\text{tw}(G)$ is the treewidth of $G$ and $\Delta(G)$ is the maximum degree of $G$[801].  [^32] The spanning tree congestion problem seeks to minimize the maximum number of detours (congestion) that pass through any edge in a spanning tree of a graph[824, 855, 801].
Proof A.5.
@@ A.1.2 Related concepts for tree-width
Related concepts include maximum order of a grid minor, maximum order of a grid-like-minor[644, 605], S-functions[1027], dimension[190], Hadwiger number[259], fractional Hadwiger number[644, 605], domino treewidth [168, 166], and the following concepts [605].
@@
  • Linear Local Tree-width: A linear variation of Local Tree-width[388, 598].
  • Simple Tree-width: Simple Tree-width holds following property.
@@
  • Circuit Tree-width [1026, 205]: Circuit Tree-width holds following property.
@@
  • Connected Tree-width [350, 926]: The minimum width of a tree-decomposition whose parts induce connected subgraphs. Connected Tree-width equals to connected tree vertex separation number[19].
@@
  • linear tree-width [561, 1185]: Linear layout width of tree-width.
  • VF-tree-width [14] :VF-tree-width is a graph width parameter defined using vertex-free tree decompositions, focusing on edges instead of vertices, useful in analyzing the complexity of cycle matroids[14].
  • Branched treewidth [464] : Branched treewidth is a graph width parameter defined by limiting the number of branching nodes in a tree decomposition path[464].
  • Radial tree-width [49] : Radial tree-width is a graph parameter that measures the smallest radial width in a tree-decomposition of a graph, where the radial width is defined by the largest radius of the subgraphs (bags) in the decomposition.
  • Loose tree-decomposition [896] : A relaxation of the tree-decomposition associated to treewidth. Related parameter is tight bramble[896]. And let $G$ be a graph and $k\geq 1$ be an integer. The following conditions are equivalent[896]:
    1. 1.
      $G$ has a loose tree-decomposition of width $k$[896];
    2. 2.
      $G$ is a minor of $T^{(k)}=T\boxdot K_{k}$ (i.e., $\mathrm{ctp}(G)\leq k$)[896];
    3. 3.
      Every tight bramble $\mathcal{B}$ of $G$ has order at most $k$[896];
    4. 4.
      The mixed search number against an agile and visible fugitive is at most $k$ (i.e., $\mathrm{avms}(G)\leq k$)[896];
    5. 5.
      The monotone mixed search number against an agile and visible fugitive is at most $k$ (i.e., $\mathrm{mavms}(G)\leq k$)[896].
  • Underlying treewidth [384] : The underlying treewidth of a graph class is the smallest treewidth of a graph such that any graph in the class is a subgraph of its product with a complete graph.
  • Twin-treewidth [991] : Twin-treewidth (ttw) is a graph parameter that generalizes treewidth. It measures the treewidth of the twin-class graph, which is obtained by contracting equivalence classes of true or false twin vertices in the original graph[991].
  • Threshold Treewidth [557] : Threshold treewidth is a refinement of treewidth that considers a graph partitioned into light and heavy vertices, with a constraint on the maximum number of heavy vertices in each bag of the tree decomposition[557].
  • supertree-width (of hypergraph) [841]: Supertree-width is a graph parameter that extends treewidth to hypergraphs by considering generalized hypertree decompositions and minimizing the maximum number of hyperedges in each decomposition[841].
  • Weak Tree-width : A method of decomposing a graph into subsets (bags) connected in a tree structure. It has looser constraints than standard tree decompositions, with relaxed connectivity and adjacency within the bags. Related concept is strong bramble[69].Let $G$ be a simple finite graph. We have
[…]
    [69]
@@ A.1.3 Application aspect for tree-width
We consider about application aspect for tree-width. Due to the versatility of tree structures in handling various types of data, tree-width has numerous applications across different fields (cf.[501]).
@@ A.2 Path-width
Path-width [763, 764, 620, 744, 645] quantifies how similar a graph is to a path by determining the minimum width of its path decomposition. An example of an obstruction to low path-width is blockage [146]. Path-width is notably applied in VLSI design [960]. Path-width is also called interval-width [769, 465].
The definition of path decomposition for a graph is as follows.
Definition A.6.
(cf.[787]) A path decomposition of a given graph $G=(V,E)$ is a sequence of subsets of $V$, $X=(X_{1},X_{2},\dots,X_{l})$, such that the following conditions hold:
  1. 1.
    $V=\bigcup_{i=1}^{l}X_{i}$ (i.e., every vertex of $G$ appears in at least one subset $X_{i}$).
  2. 2.
    For each edge $(u,v)\in E$, there exists some $i\in\{1,2,\dots,l\}$ such that both $u$ and $v$ belong to $X_{i}$.
  3. 3.
    For each $v\in V$, there exist indices $s(v),e(v)\in\{1,2,\dots,l\}$ such that $s(v)\leq e(v)$, and $v\in X_{j}$ if and only if $j\in\{s(v),s(v)+1,\dots,e(v)\}$.
The width of the path decomposition $X$ is defined as:
$\mathrm{pw}_{X}(G)=\max\{|X_{i}|\mid i=1,2,\dots,l\}-1.$
The pathwidth of $G$, denoted $\mathrm{pw}(G)$, is the minimum width over all possible path decompositions of $G$, i.e.,
$\mathrm{pw}(G)=\min\{\mathrm{pw}_{X}(G)\mid X\text{ is a path decomposition of}G\}.$
As another example, the definition of Path-width in the context of matroids is provided. For the definition of a matroid, please refer to Definition A.29 and related references.
@@ A.2.1 Hierarchy for path-width
  • For every graph, $\text{ds}(G)\leq\text{pw}(G)+1\leq(\Delta(G)+1)\text{ds}(G)$. ($\delta_{d}(G)$: the maximum degree of $G$. $\text{ds}(G)$:Domination search number)[805].
Proof A.9.
@@ A.2.2 Related concepts for path-width
  • Proper-path-width [1132, 1129, 1133, 1131, 1130, 884, 883]: Path-width that relates to mixed search game (A pursuit-evasion game on graphs where a searcher combines edge and node searches to capture a hidden intruder). The following relationship is known regarding proper-path-width.
@@
  • Connected path-width [323]: Linear layout parameter of Connected tree-width. Connected pathwidth of any graph $G$ is at most $12\times pw(G)+$ 1 [324]. Connected path-width equals to connected path vertex separation number[19].
@@
  • Radial path-width [49] : Linear layout of Radial tree-width.
  • Trellis-width [743, 462, 703, 664, 702]: Trellis-width measures the "trellis state-complexity" of a linear code, defined by the minimum dimension of intersections in a linear layout of subspaces. It is essentially equivalent to the path-width parameter.
@@ A.2.3 Application aspect for path-width
  • IoT, Network
    Applications for IoT have also been studied in the context of path-width[60, 61].
@@ A.3.1 Hierarchy for cut-width
  • Path-width is less than or equal to cut-width [788].
  • $\text{thinness}(G)\leq\text{cutwidth}(G)+1,$[875, 116]
  • Bisection width is less than or equal to cut-width[333].
@@ A.3.2 Related concepts for cut-width
  • Cyclic Cutwidth
    Cyclic Cutwidth is the minimal value of the maximum number of edges crossing any cut when the vertices are arranged on a circle[1024].
@@ A.5 Linear MIM-width
Linear MIM-width [690, 691] is the linear restriction of MIM-width.
A.5.1 Hierarchy for linear-MIM-width
We consider about Hierarchy for linear-mim-width.
Theorem A.16.
The relationships between linear-MIM-width and other graph parameters are as follows:
  • If a graph $G$ has bounded linear-MIM-width, then it also has bounded MIM-width (trivially).
  • If a graph $G$ has bounded linear-MIM-width, then it also has bounded o-mim-width and sim-width[116].
  • If a graph $G$ has bounded tree-independence-number, then it also has bounded linear-MIM-width[116].
  • If a graph $G$ has bounded path-independence-number, then it also has bounded linear-MIM-width[116].
  • If a graph $G$ has bounded simultaneous interval number, then it also has bounded linear-MIM-width[116].
  • If a graph $G$ has bounded thinness, then it also has bounded linear-MIM-width[116].
Proof A.17.
Please refer to each reference.
A.5.2 Related concepts for Linear-MIM-width
@@ A.8.1 Hierarchy for Band-width
  • If a graph $G$ has bounded Band-width, then it also has bounded bisection-width [1154].
  • If a graph $G$ has bounded Band-width, then it also has bounded Tree-width.
  • If a graph $G$ has bounded Band-width, then it also has bounded proper thinness[191].
  • If a graph $G$ has bounded band-width, then it also has bounded maximum degree, $c$-closure [782]  [^34] A graph $G$ is $c$-closed if, for all pairs of nonadjacent vertices $u$ and $v$, the size of the intersection of their neighborhoods $|N(u)\cap N(v)|$ is less than $c$. The $c$-closure of $G$ is the smallest integer $c$ for which $G$ is $c$-closed.[783, 476, 727], acyclic chromatic number, and $h$-index [425] [1154]  [^35] The $h$-index of a graph is the maximum integer $h$ such that the graph has at least $h$ vertices, each with a degree of at least $h$(cf.[426, 148])..
@@
  • For each $\Delta\geq 2$, every graph $G$ on $n$ vertices with maximum degree $\Delta(G)\leq\Delta$ has bandwidth $\text{bw}(G)\leq\frac{6n}{\log_{\Delta}\left(\dfrac{n}{\text{separator}(G)}\right)}$[199].
  • The slope number $\text{sn}(G)$ satisfies: $\text{sn}(G)\leq\frac{1}{2}\text{bw}(G)\cdot(\text{bw}(G)+1)+1.$[391]  [^36] The slope number of a graph is the minimum number of distinct edge slopes required in a straight-line drawing of the graph[1182, 700].
Proof A.22.
@@
A.8.2 Related concepts for band-width
Related parameter is Weighted band-width.
  • Edge band-width: Edge bandwidth measures the smallest maximum difference between labels of incident edges when numbering all edges of a graph; it reflects how closely edges can be labeled without large gaps[989, 835, 25].
  • Cyclic Bandwidth: Cyclic Bandwidth is the smallest possible maximum distance between adjacent vertices when arranging the vertices around a circle to minimize that maximum distance[1024].
@@ A.9 Boundary-width
Vertex-boundary-width [965] measures the minimum boundary of vertices in a discrete isoperimetric problem. Also edge-boundary-width [965] measures the minimum boundary of edges in a discrete isoperimetric problem.
@@ A.10 Carving-width
A.10.1 Hierarchy for carving-width
We consider about Hierarchy for carving-width.
Theorem A.24.
The relationships between carving-width and other graph parameters are as follows:
  • Let $G$ be a graph with maximum degree $d$. Then, the following inequality holds:
    $\frac{2}{3}\left(\text{tw}(G)+1\right)\leq\text{cng}(G)\leq d\left(\text{tw}(G)+1\right),$
    where $\text{tw}(G)$ is the tree-width of $G$ and $\text{cng}(G)$ is the carving-width of $G$.[144]
A.10.2 Related concept for Carving-width
@@
A.10.3 Application for carving-width
@@ A.11 Branch-width
Branch-width [565, 966, 1038, 177] measures the smallest width of a branch decomposition of a graph. The width is the minimum number of edges connecting subgraph pairs. This graph width parameter is known to be extendable to matroids and connectivity systems. Examples of obstructions are Tangle [565, 661], Profile [344, 345, 421, 341, 352], $k$-block [409, 430, 778], Ultrafilter [529], Maximal ideal [1231], loose tangle [966], loose tangle kit [966], Quasi-Ultrafilter [512], Weak-Ultrafilter [497], tangle kit[685], Bramble[858, 70], (k,m)-obstacle [473]. Related concepts include sphere-cut width [151] and Rooted Branch-width[704].